Grade 12th passPhysical Chemistry

In 10^3 litre sample of hard water CaSO4 and MgSO4 is present. If the elevation in boiling point is 0.000052°C. calculate the degree of hardness of hard water (Kb for H2O=0.52).

Vishnu
9 Years agoGrade 12th pass

1 Answer

Vikas TU
9 Years ago

In water testing, paper strips frequently measure hardness in parts per million (ppm), where one section for every million is characterized as one milligram of calcium carbonate (CaCO3) per liter of water. 
Molar mass of CaS04 = 110 gm 
Molar mass of mgso4 = 120gm 
Height in breaking point = 0.000052°C. Kb = 0.52. 
Level of hardness = % of cas04 = 110/10^3 *100 = 11 % 
Level of hardness = % of Mgso4= 120/10^3 *100= 12% 
Add up to degree = 12 +11= 23%